# ![[Uncaptioned image]](extracted/5527354/head_foot/header_bar.png) CryinGAN: Design and evaluation of point-cloud-based generative adversarial networks using disordered materials – application to LiScCl-LiCoO battery interfaces | Adrian Xiao Bin Yong, and Elif Ertekin |

### v1

@@ Acknowledgements
This work was supported by the Defense Advanced Research Projects Agency (DARPA) HR00112220028. We also acknowledge partial financial support from the U.S. Army CERL, awarded under Grant No. W9132T-19-2-0008. This work made use of the Illinois Campus Cluster, operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA) and supported by the University of Illinois at Urbana-Champaign. This work also utilized the Hardware-Accelerated Learning (HAL) cluster 88, supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign. This work also used Bridges-2 89 at Pittsburgh Supercomputing Center through allocation MAT220011 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@
- Zheng et al. 2023 H. Zheng, E. Sivonxay and K. Persson, MRS Spring Meeting, 2023.
@@
- Gao et al. 2021 B. Gao, R. Jalem and Y. Tateyama, ACS Applied Materials & Interfaces, 2021, 13, 11765–11773.
- He et al. 2017 S. He, K. Chen, M. Saunders, J. Li, C. Q. Cui and S. P. Jiang, Journal of The Electrochemical Society, 2017, 164, F1437.
- Wang et al. 2020 J. Wang, M. Deng, Y. Chen, X. Liu, W. Ke, D. Li, W. Dai and K. He, Materials Chemistry and Physics, 2020, 244, 122733.
- Guo et al. 2008 Y.-D. Guo, Z.-J. Yang, Q.-H. Gao and W. Dai, Physica B: Condensed Matter, 2008, 403, 2367–2371.
- Masias et al. 2019 A. Masias, N. Felten, R. Garcia-Mendez, J. Wolfenstine and J. Sakamoto, Journal of Materials Science, 2019, 54, 2585–2600.
@@
- Thanh-Tung and Tran 2020 H. Thanh-Tung and T. Tran, 2020 International Joint Conference on Neural Networks (IJCNN), 2020, pp. 1–10.
- Vignac et al. 2023 C. Vignac, N. Osman, L. Toni and P. Frossard, MiDi: Mixed Graph and 3D Denoising Diffusion for Molecule Generation, 2023.
- Samanta et al. 2020 B. Samanta, A. De, G. Jana, V. Gómez, P. Chattaraj, N. Ganguly and M. Gomez-Rodriguez, Journal of Machine Learning Research, 2020, 21, 1–33.

### v2

@@ Acknowledgements
This work was supported by the Defense Advanced Research Projects Agency (DARPA) HR00112220028. We also acknowledge partial financial support from the U.S. Army CERL, awarded under Grant No. W9132T-19-2-0008. This work made use of the Illinois Campus Cluster, operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA) and supported by the University of Illinois at Urbana-Champaign. This work also utilized the Hardware-Accelerated Learning (HAL) cluster 96, supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign. This work also used Bridges-2 97 at Pittsburgh Supercomputing Center through allocation MAT220011 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@
- Zheng et al. 2024 H. Zheng, E. Sivonxay, M. Gallant, Z. Luo, M. McDermott, P. Huck and K. A. Persson, The ab initio amorphous materials database: Empowering machine learning to decode diffusivity, 2024.
@@
- Jiao et al. 2024 R. Jiao, W. Huang, P. Lin, J. Han, P. Chen, Y. Lu and Y. Liu, Crystal Structure Prediction by Joint Equivariant Diffusion, 2024.
@@
- Kim et al. 2022 N. Kim, B. J. Blankenau, T. Su, N. H. Perry and E. Ertekin, Computational Materials Science, 2022, 202, 110969.
- Su et al. 2024 T. Su, B. J. Blankenau, N. Kim, J. A. Krogstad and E. Ertekin, Acta Materialia, 2024, 276, 120088.
- Cowley 1965 J. M. Cowley, Physical Review, 1965, 138, A1384–A1389.
- Bartók et al. 2018 A. P. Bartók, J. Kermode, N. Bernstein and G. Csányi, Physical Review X, 2018, 8, 041048.
@@
- Jain et al. 2013 A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder and K. A. Persson, APL Materials, 2013, 1, 1–11.
- Wang 2024 P. Wang, imagen-pytorch, 2024, https://github.com/lucidrains/imagen-pytorch.
- Ho et al. 2022 J. Ho, T. Salimans, A. Gritsenko, W. Chan, M. Norouzi and D. J. Fleet, Video Diffusion Models, 2022.
@@
- Dhariwal and Nichol 2021 P. Dhariwal and A. Nichol, Advances in Neural Information Processing Systems, 2021, pp. 8780–8794.
- Müller-Franzes et al. 2023 G. Müller-Franzes, J. M. Niehues, F. Khader, S. T. Arasteh, C. Haarburger, C. Kuhl, T. Wang, T. Han, T. Nolte, S. Nebelung, J. N. Kather and D. Truhn, Scientific Reports, 2023, 13, 12098.
- Davies et al. 2016 D. Davies, K. Butler, A. Jackson, A. Morris, J. Frost, J. Skelton and A. Walsh, Chem, 2016, 1, 617–627.
@@
- Sanchez et al. 1984 J. M. Sanchez, F. Ducastelle and D. Gratias, Physica A: Statistical Mechanics and its Applications, 1984, 128, 334–350.
- Metropolis et al. 1953 N. Metropolis, A. W. Rosenbluth, M. N. Rosenbluth, A. H. Teller and E. Teller, The Journal of Chemical Physics, 1953, 21, 1087–1092.
- Kawasaki 1966 K. Kawasaki, Physical Review, 1966, 145, 224.
- Saharia et al. 2022 C. Saharia, W. Chan, S. Saxena, L. Li, J. Whang, E. Denton, S. K. S. Ghasemipour, B. K. Ayan, S. S. Mahdavi, R. G. Lopes, T. Salimans, J. Ho, D. J. Fleet and M. Norouzi, Photorealistic Text-to-Image Diffusion Models with Deep Language Understanding, 2022.
- Morgan 2021 B. J. Morgan, vasppy, 2021, https://pypi.org/project/vasppy/.